Hamid left his home and travel 10Km towards Sarafah bazaar with the velocity of 10m/s.

He then moves towards Empress Market with the same velocity and travel 5Km. After that, he returned to his home on the resultant path. Suppose he travelled with constant velocity of 10m/s.

Sarafah bazaar is in west of Hamid’s home, and Empress market south of Sarafah bazaar

a) Find the time it will take him to reach his home.

b) Solve part (a), considering that he has doubled his speed in his return journey.


1
Expert's answer
2020-11-23T10:26:39-0500

As per the given question,

Distance between Hamid's home and Sarafah bazaar "(d_1)=10KM=10000m"

Hamid's velocity "(v)=10m\/s"

Distance between Sarafah bazaar and Empress market "(d_2)=5KM = 5000m"



Hence, the net displacement between Hamid's home and Empress market "d=\sqrt{d_1^2+d_2^2}"

Now, substituting the values, "d=\sqrt{10^2+5^2}=\sqrt{125}=11.18KM"

"d=11180m"

Hence time taken to reach the home "t=\frac{11180}{10} =1118sec"

if the returning speed gets double then time taken by Hamid to reach home"=\frac{1118}{20}=55.9 sec"